OpenStudy (anonymous):

If f(x)=10^x and 10^1.04 is 10.96, which is closest to f'(1)? a)0.24 b)0.92 c)0.96 d)10.5 e)24

OpenStudy (anonymous):

@Michele_Laino

OpenStudy (michele_laino):

we have to find the first derivative of 10^x

OpenStudy (anonymous):

that is 10^x*ln10 right?

OpenStudy (michele_laino):

that's right!

OpenStudy (michele_laino):

so what is: \[f'\left( 1 \right) = {10^1} \times \ln 10 = ...?\]

OpenStudy (anonymous):

its 10ln10, I cannot use a calculator for this

OpenStudy (michele_laino):

ok!

OpenStudy (anonymous):

I believe the question has something to do with local linearization but I'm not sure

OpenStudy (freckles):

\[f(x) \approx f'(1.04)(x-1.04)+f(1.04)\]

OpenStudy (anonymous):

I don't know f'(1.04)

OpenStudy (freckles):

you can find f' from f by differentiating f

OpenStudy (anonymous):

that gives me 10.96ln10

OpenStudy (michele_laino):

we can linearize the function 10^x around the point x=1.04

OpenStudy (anonymous):

ok

OpenStudy (freckles):

the only bad thing is they tell us what they want us to use for 10^(1.04) but they don't tell us what they want us to use for ln(10)

OpenStudy (anonymous):

thats my main problem

OpenStudy (freckles):

and you aren't allowed to use a calculator to figure out ln(10), right?

OpenStudy (michele_laino):

yes!

OpenStudy (anonymous):

unfortunately no

OpenStudy (michele_laino):

do you know wath is log_10(e) ?

OpenStudy (anonymous):

nope

OpenStudy (freckles):

\[\ln(10)=y \\ \text{ equivalently this is } e^y=10 \\ \text{ we know } e \approx 2.7 \\ 2.7^y=10 \\ \text{ this equation would give a } y \text{ close \to this other equation's y } \\3^y=10\] y is like 2.something

OpenStudy (freckles):

let's use the approximate that y is 2.1 and see what we get at the end

OpenStudy (freckles):

we will use ln(10) is approximately 2.1

OpenStudy (anonymous):

ok sounds good

OpenStudy (freckles):

\[f(x) \approx f'(1.04)(x-1.04)+f(1.04) \\ f'(x) \approx 10.96 \cdot 2.1(x-1.04)+10.96 \]

OpenStudy (anonymous):

and then plug 1 into that right?

OpenStudy (freckles):

yah

OpenStudy (anonymous):

ok thank you

OpenStudy (michele_laino):

another way is : since we have: e < 10 <e^(5/2) then we can write: 1<ln(10)<5/2 and finally 10.96 < f '(1) <25

OpenStudy (michele_laino):

10 < f '(1) <25

OpenStudy (freckles):

I think something is wrong above

OpenStudy (freckles):

not with what @Michele_Laino said

OpenStudy (freckles):

I'm so dumb

OpenStudy (freckles):

I accidentally changed f(x) to f'(x)

OpenStudy (freckles):

\[f(x) \approx f'(1.04)(x-1.04)+f(1.04) \\ f(x) \approx 10.96 \cdot 2.1(x-1.04)+10.96\] this is what was suppose to say but we were looking for f'(1) not f(1)

OpenStudy (anonymous):

oh right

OpenStudy (freckles):

lol we should have just found f' given f

OpenStudy (freckles):

\[f'(x)=\ln(10) \cdot 10^x \\ f'(1)=\ln(10) \cdot 10^{1}\]

OpenStudy (freckles):

I honestly don't know why they gave you 10^(1.04)

OpenStudy (freckles):

\[f'(1)=2.1 \cdot 10\]

OpenStudy (freckles):

which is closer to that 23 choice

OpenStudy (anonymous):

lol i guess it was easier than i thought

OpenStudy (anonymous):

brute force seems to work

OpenStudy (freckles):

maybe it was a trick

OpenStudy (freckles):

like I bet my students will try to use that liberalization method if I say this

OpenStudy (anonymous):

you should try it on them and see what happens XD

OpenStudy (freckles):

you know what

OpenStudy (freckles):

actually we still could use that lineazation method

OpenStudy (freckles):

oops one sec

OpenStudy (freckles):

\[f(x) \approx f'(a)(x-a)+f(a) \\ f(x) \approx f'(1)(x-1)+f(1) \\ \text{ we want \to solve for } f'(1) \\ f(x)-f(1) \approx f'(1)(x-1) \\ \frac{f(x)-f(1)}{x-1} \approx f'(1)\]

OpenStudy (freckles):

and we were given f(x) for when x=1.04

OpenStudy (freckles):

\[f'(1) \approx \frac{f(x)-f(1)}{x-1} \text{ @ x=1.04 } \\ f'(1) \approx \frac{10^{1.04}-10^{1}}{1.04-1} \approx \frac{10.96-10}{1.04-1}\]

OpenStudy (freckles):

how sick is that! :)

OpenStudy (anonymous):

ah thats a lot better

OpenStudy (anonymous):

very satisfying

OpenStudy (freckles):

I am actually more satisfied with that too

OpenStudy (anonymous):

i think thats what the question makers were going for

OpenStudy (freckles):

yep yep me too

OpenStudy (freckles):

it wasn't a trick after all

OpenStudy (anonymous):

ok well thanks a ton for helping!

OpenStudy (freckles):

np